The sum of two numbers is 18 and their difference is 4. What are the two numbers?
Answer by scott8148(6628)

let x=number, y=(other number) ... x+y=18 ... x-y=4

adding the two equations eliminates the y and gives 2x=22, so x=11

11+y=18 ... y=7